# DTG : Diffusion-based Trajectory Generation for Mapless Global Navigation

### 14 Mar 2024 v1

@@ I Introduction
Computing Traversable Trajectories for Robotic Navigation: In complex outdoor scenarios, traversability analysis is critical for safe outdoor navigation because outdoor environments contain various challenging terrains, plants, buildings, trees, etc. [liang2022adaptiveon, weerakoon2023graspe, xiao2021learning, karnan2022vi, pokhrel2024cahsor, jardali2024autonomous]. Traditionally, perception and planning are decoupled as two different tasks [canny1988complexity, manocha1992algebraic, lavalle2006planning]: first, use perception models to detect traversable areas for the robot, taking its dynamic constraints into account; second, apply traversability constraints to a planner and generate a feasible trajectory to move the robot. However, the traversable maps created by these approaches may not be the most effective representation of the environment. For example, in learning-based approaches [liang2022adaptiveon, liang2021crowd, sridhar2023nomad, shah2023vint], the high-dimensional observations are usually processed into a vector to encode the traversability information. Furthermore, for mapless navigation with occlusions, estimating traversable areas behind occluding objects can be challenging. Therefore, effectively modeling the complex traversability information in outdoor scenarios is important.
Optimal Trajectory Generation Towards Designated Goals. Beyond ensuring traversability, navigation problems are often formulated as optimization problems [liang2021vo, liang2022adaptiveon, malyuta2022convex], aiming to compute optimal trajectories under constraints. Some of the common optimality criteria include minimizing the running time or the traveling distance [liang2021crowd, malyuta2022convex, gasparetto2015path, sanchez2021path]. However, in mapless navigation, if the goal is far away, the absence of an exact map complicates the evaluation of the remaining travel distance, e.g., whether traversing through a wider or narrower passage will result in shorter travel distance [liang2023mtg, wu2022image, jang2021hindsight].
@@
    A Novel Conditional RNN (CRNN) Model for The Diffusion Model: To make diffusion models run in real time for trajectory generation, we propose CRNN, which takes the environment information as conditions and generates trajectories in real time for global navigation.
@@
    Adaptive Training to Enhance Traversability: To enhance the traversability of the generated trajectories, we propose a new method to adaptively add traversability loss to train the diffusion-based trajectory generation. The traversability loss is adaptively applied to different diffusion steps according to the historic loss of the steps.
@@ II Prior Work and Background
Outdoor Navigation: Navigating robots in outdoor environments presents significant challenges due to occlusions, weather conditions, construction sites, and diverse and complex terrains like grass, bushes, mud, and sharp elevation changes [weerakoon2023graspe, xiao2021learning, karnan2022vi, pokhrel2024cahsor, datar2023toward, canny1988complexity, manocha1992algebraic]. Various strategies [liang2022adaptiveon, weerakoon2023graspe, xu2023benchmarking, xu2021applr, xu2021machine] have been proposed to address these issues. Motion planning techniques aim for stable and safe robot movement across different terrains [liang2022adaptiveon, weerakoon2023graspe], using adaptive approach [liang2022adaptiveon] and reinforcement learning [xu2023benchmarking, xu2021applr, xu2021machine] to train neural networks for waypoint generation. Global planning requires a comprehensive cost map for path planning [fusic2021optimal, li2021openstreetmap] and accurate robot localization [akai2017robust, lowry2015visual] to follow the paths. However, those map-based navigation approaches can be computationally expensive and require a significant amount of overhead to maintain the maps. To solve this issue, instead of maintaining a comprehensive map, Shah et al. [shah2023vint], Sridhar et al. [sridhar2023nomad], and Hirose et al. [hirose2023exaug] propose generating topological maps and using images as subgoals for navigation, but those approaches still require initial runs in the environment to gather subgoal images. Mapless navigation techniques are used to navigate without relying on maps. Giovannangeli et al.[giovannangeli2006robust] and Liang et al. [liang2022adaptiveon] generate actions in a mapless manner, but they focuse on local planning without addressing long-distance navigation. MTG [liang2023mtg] represents a step forward by providing long-distance navigation trajectories, yet it doesn’t optimize for the best path. In contrast, we propose a novel approach, DTG, which generates traversable trajectories with short travel distance towards a distant goal in large-scale outdoor settings without a map.
@@
Traversability Analysis: Traversability analysis plays a critical role in robot navigation, distinguishing between navigable and non-navigable areas within an environment. This task is often handled by a separation of perception and planning [fan2021step, gasparino2022wayfast, jonasfast, sikand2022visual], utilizing various sensors to assess the terrain. Cameras (RGB or RGB-D) are commonly employed to analyze the terrain [hosseinpoor2021traversability, jonasfast], enabling segmentation and the creation of cost maps for navigation. Similarly, Lidar sensors are used for their ability to generate elevation maps through geometric information [xue2023traversability, Fankhauser2018ProbabilisticTerrainMapping]. However, reliance on cost or elevation maps, while visually intuitive for humans, poses computational costs for robots due to the extra processing and encoding of the maps. To address this, end-to-end learning approaches [xiao2022motion, liang2022adaptiveon, liang2023mtg] have been developed to encode environmental information directly into neural networks, thereby streamlining the navigation process. In our work, we also imply an end-to-end learning approach to encode the observation information.
@@
Trajectory Generation: Trajectory generation in outdoor scenarios varies significantly for different applications. Some methods are designed specifically for autonomous driving [varadarajan2022multipath, nayakanti2023wayformer, shi2022motion] with LSTM [altche2017lstm] or Gaussian Mixture Models [varadarajan2022multipath] to predict trajectories based on historical movements. Smaller robots’ trajectory generation strategies [kahn2021land, shah2022viking, shah2023vint] leverage Bayesian-based methods [liang2023mtg, yang2020online], GANs (Generative Adversarial Networks) [gupta2018social, 10323465] etc to compute feasible paths through complex environments. This approach specifically addresses the challenges of navigating small robots in a large-scale mapless environment. Existing methods, such as ViNT [shah2023vint] and NoMaD [sridhar2023nomad], rely on comparing current images to pre-recorded subgoal images to navigate, but they cannot recognize the perceived images with significant differences from the subgoal images or in completely unknown environments, so sophisticated choices of subgoals are necessary [shah2023vint]. Moreover, they do not consider the optimality of the trajectories in terms of travel distance to the goals and require prior knowledge (subgoal images) of the environment, limiting their applicability to unknown or dynamically changing areas. MTG [liang2023mtg] uses CVAE [cvae] to generate trajectories in traversable areas, but it doesn’t consider the optimality of the trajectories. Diffusion models [ho2020denoising, song2020denoising] have been used for robotics applications, such as picking and pushing objects [chi2023diffusion] and navigation [sridhar2023nomad]. In our approach, we propose a novel diffusion-based mapless trajectory generator to generate optimal trajectories with short travel distance to goals and also train the generation with traversabilities metrics.
@@ III Approach
Figure 2: DTG Architecture: DTG has two models: Perception Encoder and Diffusion Model. Perception Encoder encodes the observation information, $\mathbf{o}=\left\{o_{l},o_{v},g\right\}$, to the condition vector, $\mathbf{c}$. The diffusion Model takes a Gaussian distribution to a trajectory $\hat{\bm{\tau}}$ under the condition $\mathbf{c}$.
@@ III-A Problem Definition
The problem we are solving is outdoor mapless global navigation. Given a very distant goal $g\in\mathcal{O}_{g}$, which is assumed to be mostly out of the vision of the robot in a large-scale environment, our model generates trajectories in the robot’s traversable areas while trying to minimize the travel distance to the goal. We assume the trajectory generator ignores small and dynamic obstacles, which can be handled by the local planners.
Perceptual sensors used in our approach include a 3D LiDAR and the robot’s odometer. We utilize $C_{l}$ consecutive frames of LiDAR perception, $\mathcal{O}_{l}$, to capture the static and dynamic information of the environment in the robot’s vicinity. The odometer provides the latest $C_{v}$ consecutive frames of the robot’s velocities, $\mathcal{O}_{v}$, to encode the robot’s dynamic status. In this global navigation task, the goal denotes as $g\in\mathcal{O}_{g}$. Thus, the observation contains $\mathcal{O}=\left\{\mathcal{O}_{l},\mathcal{O}_{v},\mathcal{O}_{g}\right\}$. Given the observation $\mathcal{O}$, our trajectory generator $DTG_{\theta}$ generates a trajectory $\bm{\tau}=\{\mathbf{w}_{1},...,\mathbf{w}_{M}\}$ that satisfies both traversability constraints and has minimum travel distance to the goal $g$. Here, $\mathbf{w}_{m}=\{x_{m},y_{m}\}$ represents the waypoints in the generated trajectory, which contains M waypoints in total starting from the robot. We define the travel distance of the trajectory $\bm{\tau}$ as the length of the shortest path from the last waypoint $\mathbf{w}_{M}$ of $\bm{\tau}$ to the target $g$. For each observation $o\in\mathcal{O}$, we have $DTG_{\theta}(\mathbf{o})=\bm{\tau}$. The problem can be formulated as the following equation:
@@
where $h(\cdot,\cdot)$ represents the travel-distance function between the last waypoint $\mathbf{w}_{M}$ and goal $g$. $\mathcal{A}$ is the traversable area around the robot and $\tilde{\mathcal{A}}$ represents non-traversable areas. $\beta$ is a hyperparameter and $f(\cdot,\cdot)$ calculates the traversability of the trajectory $\bm{\tau}$, which is the portion of non-traversable areas covered by the trajectory $\bm{\tau}$. Therefore, the function achieves two targets: 1. optimize the travel distance; and 2. satisfy traversability constraints. The inference of our approach only takes $\mathbf{o}\in\mathcal{O}$ as input, but to train the models, we require $\mathcal{A}$ to calculate the traversability ground truth.
For training, DTG uses a similar traversability map as in MTG [liang2023mtg], where the off-road areas and buildings are not traversable but side walks, pavements and motor drive road are traversable areas. The travel-distance function $h(\cdot,\cdot)$ can be handled by the A* algorithm, starting from the last waypoint $\mathbf{w}_{M}$ to the goal $g$, and the travel distance is the length of the calculated A* path. The non-traversable area $\tilde{\mathcal{A}}$ can also be directly extracted from the traversability map.
@@
where $\bm{\tau}_{gt}$ is the ground truth trajectory with the shortest travel distance to the goal $g$ and in traversable areas. $d(\cdot,\cdot)$ calculates the distance between the generated trajectory and the ground truth trajectory $\bm{\tau}_{gt}$. Since the ground truth and inputs are all given, this problem can be handled by a supervised learning method.
@@ III-B Architecture
Figure 2 shows the end-to-end architecture of our approach, DTG. There are two models in the pipeline: Perception Encoder $P_{\theta}(\cdot)$ and Diffusion Model $D_{\theta}(\cdot)$. Here we denote all model parameters as $\theta$. The detail of the layer configurations is in the Appendix VI.
@@ III-B1 Perception Encoder
where $o_{l}\in\mathcal{O}_{l}$, $o_{v}\in\mathcal{O}_{v}$ and $g\in\mathcal{O}_{g}$. $p^{c}_{\theta}(\cdot)$ represents the model PointCNN [hua2018pointwise]. $p^{v}_{\theta}(\cdot)$ is a sequence of Linear layers to process the robot’s dynamic status with historic velocities. As shown in Figure 2, Encoder layer takes the concatenated embeddings from $p^{c}_{\theta}(o_{l}),p^{v}_{\theta}(o_{v})$ and $g$ and encode the embeddings to a vector, $\mathbf{c}$, as the condition of the Diffusion model. The Encoder, $p^{e}_{\theta}(\cdot)$, composes a sequence of Linear layers, as well.
@@ III-B2 Diffusion Model
Diffusion model generates high-quality data by progressively denoising a Guassian noise to some target data [ho2020denoising, song2020denoising], and it shows promising capability in robotics tasks [chi2023diffusion], as well. In our approach, the diffusion model takes the conditional vector $\mathbf{c}$ from the Perception Encoder for each observation and denoise a Gaussian distribution to a trajectory $\hat{\bm{\tau}}$. In recent years, multiple diffusion training strategies have been proposed, including DDPM [ho2020denoising], DDIM [song2020denoising], and VDM [vdm]. Because VDM [vdm] shows higher quality in data generation and better convergence than others, in this approach, we use the VDM [vdm] strategy to generate trajectories. In diffusion process, neural networks (diffusion cells) are trained to learn noise or noisy data, and these two procedures are mathematically equivalent. Because of faster convergence and smaller converged loss, we use the diffusion cells to predict trajectories in our approach, DTG. Our diffusion generator contains N steps. As shown in Figure 2, from the final output ($\mathbf{x}_{0}$) to the Gaussian noise ($\mathbf{x}_{N}$), noise is introduced in each step, resulting in a noisy trajectory denoted as $\tilde{\mathbf{x}}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\alpha_{t}$ is the noise ratio at step $t$, which ranges from 1 to N. $\epsilon$ represents the noise itself. The predicted trajectory in step $t$ is $\hat{\mathbf{x}}_{t-1}=R_{\theta}(\tilde{\mathbf{x}}_{t},t)$, where $R_{\theta}(\cdot)$ is a diffusion cell. Because the U-Net [ho2020denoising, chi2023diffusion] is very computationally enhansive, we propose a novel diffusion cell to reduce the computational cost in Figure 2. Since we need the generated trajectories to always remain within traversable areas, we require the diffusion cell to also integrate environmental information. Inspired by [chi2023diffusion], we propose a Conditional RNN (CRNN), right side of Figure 2, that takes environmental information vector $\mathbf{c}$ and the step number $t$ as conditions for the generator, denoted by $R_{\theta}(\tilde{\mathbf{x}}_{t-1},t,\mathbf{c})$. Thus, we have the CRNN cell:
@@
where $d^{h}_{\theta}(\cdot)$ calculates the hidden vector $\mathbf{h}$ for GRU cells. $d^{p}(\cdot)$ represents Sinusoidal positional embedding. $f^{1}_{\theta}(\cdot)$ and $f^{2}_{\theta}(\cdot)$ are all Linear layers. $k\in\left\{1,...,K\right\}$ represents the steps in the CRNN model, and each step is $r^{k}_{\theta}(\cdot)$. $d^{g}_{\theta}(\cdot)$ is GRU cell and $\mathbf{h}$ is the hidden vector for the GRU cell. We also compare with U-Net [chi2023diffusion] as the conditional encoder and show the results in Table I and demonstrate that we have comparable results, in terms of traversability and travel distance, but significantly less computational cost.

### 25 Mar 2024 v2

@@ I Introduction
Computing Traversable Trajectories for Robotic Navigation: In complex outdoor scenarios, traversability analysis is critical for safe outdoor navigation because outdoor environments contain various challenging terrains, plants, buildings, trees, etc. [liang2022adaptiveon, weerakoon2023graspe, xiao2021learning, karnan2022vi, pokhrel2024cahsor]. Traditionally, perception and planning are decoupled as two different tasks [canny1988complexity, manocha1992algebraic, lavalle2006planning], perception to detect traversable areas and planning to generate waypoints. However, the traversable maps created by these approaches may not be the most effective representation of the environment. For example, in learning-based approaches [liang2021crowd, sridhar2023nomad, shah2023vint], the high-dimensional observations are usually processed into a vector to encode the traversability information. Furthermore, for mapless navigation with occlusions, estimating traversable areas behind occluding objects can be challenging. Therefore, effectively modeling the complex traversability information in outdoor scenarios is important.
Optimal Trajectory Generation Towards Designated Goals. Beyond ensuring traversability, navigation problems are often formulated as optimization problems [liang2021vo, liang2022adaptiveon, malyuta2022convex], aiming to compute optimal trajectories under constraints. Some of the common optimality criteria include minimizing the running time or the traveling distance [liang2021crowd, malyuta2022convex, gasparetto2015path, sanchez2021path]. However, in mapless navigation, if the goal is far away, the absence of an exact map complicates the evaluation of the remaining travel distance, e.g., whether traversing through a wider or narrower passage will result in shorter travel distance [wu2022image, jang2021hindsight].
@@
    A Novel Conditional RNN (CRNN) Model for The Diffusion Model: To make diffusion models run in real time for navigation, we propose CRNN, which takes the environment information as conditions and generates trajectories in real time for global navigation.
@@
    Adaptive Training to Enhance Traversability: To enhance the traversability of the generated trajectories, we propose a new method to adaptively apply traversability loss to different diffusion steps according to the historic loss of the steps.
@@ II Prior Work and Background
Outdoor Navigation: Navigating robots in outdoor environments presents significant challenges due to occlusions, weather conditions, construction sites, and diverse and complex terrains like grass, bushes, mud, and sharp elevation changes [weerakoon2023graspe, xiao2021learning, karnan2022vi, pokhrel2024cahsor, canny1988complexity, manocha1992algebraic]. Various strategies [liang2022adaptiveon, weerakoon2023graspe, xu2021applr, xu2021machine] have been proposed to address these issues. Motion planning techniques aim for stable and safe robot movement across different terrains [liang2022adaptiveon, weerakoon2023graspe], using adaptive approach [liang2022adaptiveon] and reinforcement learning [xu2021applr, xu2021machine] to train neural networks for waypoint generation. Global planning requires a comprehensive cost map for path planning [fusic2021optimal, li2021openstreetmap] and accurate robot localization [akai2017robust, lowry2015visual] to follow the paths. However, those map-based navigation approaches can be computationally expensive and require a significant amount of overhead to maintain the maps. To solve this issue, instead of maintaining a comprehensive map, Sridhar et al. [shah2023vint, sridhar2023nomad] and Hirose et al. [hirose2023exaug] propose generating topological maps and using images as subgoals for navigation, but those approaches still require initial runs in the environment to gather subgoal images. Mapless navigation techniques are used to navigate without relying on maps. Giovannangeli et al.[giovannangeli2006robust] and Liang et al. [liang2022adaptiveon] generate actions in a mapless manner, but they focus on local planning without addressing long-distance navigation. MTG [liang2023mtg] represents a step forward by providing long-distance navigation trajectories, yet it doesn’t optimize for the best path. In contrast, we propose a novel approach, DTG, which generates traversable trajectories with short travel distances towards a distant goal in large-scale outdoor settings without a map.
@@
Traversability Analysis: Traversability analysis plays a critical role in robot navigation, distinguishing between navigable and non-navigable areas within an environment. This task is often handled by separating perception and planning [fan2021step, gasparino2022wayfast, jonasfast], utilizing various sensors to assess the terrain. Cameras (RGB or RGB-D) are commonly employed to analyze the terrain [hosseinpoor2021traversability, jonasfast], enabling segmentation and the creation of cost maps for navigation. Similarly, Lidar sensors are used for their ability to generate elevation maps through geometric information [xue2023traversability, Fankhauser2018ProbabilisticTerrainMapping]. However, reliance on cost or elevation maps, while visually intuitive for humans, poses computational costs for robots due to the extra processing and encoding of the maps. To address this, end-to-end learning approaches [xiao2022motion, liang2022adaptiveon, liang2023mtg] have been developed to encode environmental information directly into neural networks, thereby streamlining the navigation process. In our work, we also apply an end-to-end learning approach to encode the observation information.
@@
Trajectory Generation: Trajectory generation in outdoor scenarios varies significantly for different applications. Some methods are designed specifically for autonomous driving [varadarajan2022multipath, nayakanti2023wayformer, shi2022motion] with LSTM [altche2017lstm] or Gaussian Mixture Models [varadarajan2022multipath] to predict trajectories based on historical movements. Smaller robots’ trajectory generation strategies [kahn2021land, shah2022viking, shah2023vint] leverage Bayesian-based methods [liang2023mtg, yang2020online], GANs (Generative Adversarial Networks) [gupta2018social, 10323465] etc to compute feasible paths through complex environments. Existing methods for small robots global navigation, such as ViNT [shah2023vint] and NoMaD [sridhar2023nomad], rely on comparing current images to pre-recorded subgoal images to navigate, but they cannot recognize the perceived images with significant differences from the subgoal images or in completely unknown environments, so sophisticated choices of subgoals are necessary [shah2023vint]. Moreover, they do not consider the optimality of the trajectories in terms of travel distance to the goals and require prior knowledge (subgoal images) of the environment, limiting their applicability to unknown or dynamically changing areas. MTG [liang2023mtg] uses CVAE [cvae] to generate trajectories in traversable areas, but it doesn’t consider the optimality of the trajectories. Diffusion models [ho2020denoising, song2020denoising] have been used for robotics applications, such as picking and pushing objects [chi2023diffusion] and navigation [sridhar2023nomad]. In our approach, we propose a novel diffusion-based mapless trajectory generator to generate optimal trajectories with short travel distances to goals and train the generation with traversabilities metrics.
@@ III Approach
Figure 2: DTG Architecture: DTG has two models: Perception Encoder and Diffusion Model. Perception Encoder encodes the observation information, $\mathbf{o}=\left\{o_{l},o_{v},g\right\}$, to the condition vector, $\mathbf{c}$. The Diffusion Model takes a Gaussian distribution to generate a trajectory $\hat{\bm{\tau}}$ under the condition $\mathbf{c}$.
@@ III-A Problem Definition
The problem we are solving is outdoor mapless global navigation. Given a distant goal $g\in\mathcal{O}_{g}$ in a large-scale environment, our model generates trajectories in the robot’s traversable areas while trying to minimize the travel distance to the goal. We assume the trajectory generator ignores small and dynamic obstacles, which can be handled by the local planners.
Perceptual sensors used in our approach include a 3D LiDAR and the robot’s odometer. We utilize $C_{l}$ consecutive frames of LiDAR perception, $\mathcal{O}_{l}$, to capture the static and dynamic information of the environment in the robot’s vicinity. The odometer provides the latest $C_{v}$ consecutive frames of the robot’s velocities, $\mathcal{O}_{v}$, to encode the robot’s dynamic status. In this global navigation task, the goal is denoted as $g\in\mathcal{O}_{g}$. Thus, the observation contains $\mathcal{O}=\left\{\mathcal{O}_{l},\mathcal{O}_{v},\mathcal{O}_{g}\right\}$. Given the observation $\mathcal{O}$, our trajectory generator $DTG_{\theta}$ generates a trajectory $\bm{\tau}=\{\mathbf{w}_{1},...,\mathbf{w}_{M}\}$ that both satisfies traversability constraints and has minimum travel distance to the goal $g$. Here, $\mathbf{w}_{m}=\{x_{m},y_{m}\}$ represents the waypoints in the generated trajectory, which contains M waypoints in total starting from the robot. We define the travel distance of the trajectory $\bm{\tau}$ as the length of the shortest path from the last waypoint $\mathbf{w}_{M}$ of $\bm{\tau}$ to the target $g$. For each observation $o\in\mathcal{O}$, we have $DTG_{\theta}(\mathbf{o})=\bm{\tau}$. The problem can be formulated as the following equation:
@@
where $h(\cdot,\cdot)$ represents the travel-distance function between the last waypoint $\mathbf{w}_{M}$ and the goal $g$. $\mathcal{A}$ is the traversable area around the robot and $\tilde{\mathcal{A}}$ represents non-traversable areas. $\beta$ is a hyperparameter and $f(\cdot,\cdot)$ calculates the traversability of the trajectory $\bm{\tau}$, which is the portion of non-traversable areas covered by the trajectory $\bm{\tau}$. Therefore, the function achieves two targets: 1. Optimize the travel distance; and 2. Satisfy traversability constraints. The inference of our approach only takes $\mathbf{o}\in\mathcal{O}$ as input, but to train the models, we require $\mathcal{A}$ to calculate the traversability ground truth.
For training, DTG uses a similar traversability map as in MTG [liang2023mtg], where the off-road areas and buildings are not traversable, but sidewalks, pavements, and drivable roads are traversable areas. The travel-distance function $h(\cdot,\cdot)$ can be handled by the A* algorithm, starting from the last waypoint $\mathbf{w}_{M}$ to the goal $g$. The travel distance is the length of the calculated A* path. The non-traversable area $\tilde{\mathcal{A}}$ can also be directly extracted from the traversability map.
@@
where $\bm{\tau}_{gt}$ is the ground truth trajectory with the shortest travel distance to the goal $g$ and lies in traversable areas. $d(\cdot,\cdot)$ calculates the distance between the generated trajectory and the ground truth trajectory $\bm{\tau}_{gt}$. Since the ground truth and inputs are all given, this problem can be handled by a supervised learning method.
@@ III-B Architecture
Figure 2 shows the end-to-end architecture of our approach, DTG. There are two models in the pipeline: Perception Encoder $P_{\theta}(\cdot)$ and Diffusion Model $D_{\theta}(\cdot)$. Here we denote all model parameters as $\theta$. Details about the layer configurations are in Appendix VI[supplements].
@@ III-B1 Perception Encoder
where $o_{l}\in\mathcal{O}_{l}$, $o_{v}\in\mathcal{O}_{v}$ and $g\in\mathcal{O}_{g}$. $p^{c}_{\theta}(\cdot)$ represents the model PointCNN [hua2018pointwise]. $p^{v}_{\theta}(\cdot)$ is a sequence of Linear layers to process the robot’s dynamic status with historic velocities. As shown in Figure 2, the Encoder layer takes the concatenated embeddings from $p^{c}_{\theta}(o_{l}),p^{v}_{\theta}(o_{v})$, and $g$ and encodes the embeddings to a vector, $\mathbf{c}$, as the condition of the Diffusion model. The Encoder, $p^{e}_{\theta}(\cdot)$, also composes a sequence of Linear layers.
@@ III-B2 Diffusion Model
The Diffusion model generates high-quality data by progressively denoising a Gaussian noise to some target data [ho2020denoising, song2020denoising], and it shows promising capabilities in robotics tasks [chi2023diffusion]. In our approach, the diffusion model takes the conditional vector $\mathbf{c}$ from the Perception Encoder for each observation and denoise a Gaussian distribution to a trajectory $\hat{\bm{\tau}}$. In recent years, multiple diffusion training strategies have been proposed, including DDPM [ho2020denoising], DDIM [song2020denoising], and VDM [vdm]. Because VDM [vdm] shows higher quality in data generation and better convergence than others, in this approach, we use the VDM [vdm] strategy to generate trajectories. In the diffusion process, neural networks (diffusion cells) are trained to learn noise or noisy data, and these two procedures are mathematically equivalent. Because of faster convergence and smaller converged loss, we use the diffusion cells to predict trajectories in our approach, DTG. Our diffusion generator contains N steps. As shown in Figure 2, from the final output ($\mathbf{x}_{0}$) to the Gaussian noise ($\mathbf{x}_{N}$), noise is introduced in each step, resulting in a noisy trajectory denoted as $\tilde{\mathbf{x}}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\alpha_{t}$ is the noise ratio at step $t$, which ranges from 1 to N. $\epsilon$ represents the noise itself. The predicted trajectory in step $t$ is $\hat{\mathbf{x}}_{t-1}=R_{\theta}(\tilde{\mathbf{x}}_{t},t)$, where $R_{\theta}(\cdot)$ is a diffusion cell. Because the U-Net [ho2020denoising, chi2023diffusion] is very computationally enhancive, we propose a novel diffusion cell to reduce the computational cost in Figure 2. Since we need the generated trajectories to remain always within traversable areas, we require the diffusion cell to also integrate environmental information. Inspired by [chi2023diffusion], we propose a Conditional RNN (CRNN), shown on the right side of Figure 2, that takes environmental information vector $\mathbf{c}$ and the step number $t$ as conditions for the generator, denoted by $R_{\theta}(\tilde{\mathbf{x}}_{t-1},t,\mathbf{c})$. Thus, we have the CRNN cell:
@@
where $d^{h}_{\theta}(\cdot)$ calculates the hidden vector $\mathbf{h}$ for GRU cells. $d^{p}(\cdot)$ represents Sinusoidal positional embedding. $f^{1}_{\theta}(\cdot)$ and $f^{2}_{\theta}(\cdot)$ are all Linear layers. $k\in\left\{1,...,K\right\}$ represents the steps in the CRNN model. Each step is $r^{k}_{\theta}(\cdot)$. $d^{g}_{\theta}(\cdot)$ is the GRU cell and $\mathbf{h}$ is the hidden vector for the GRU cell. We also compare with U-Net [chi2023diffusion] as the conditional encoder and show the results in Table I demonstrating that we have comparable results, in terms of traversability and travel distance, but significantly less computational cost.